OpenStudy (anonymous):

A chemist prepares a solution of potassium permanganate (KMnO4) by measuring out 25.7 micro moles of potassium permanganate into a volumetric flask and filling the flask to the mark with water.Calculate the concentration in mol/L of the chemist's potassium permanganate solution. Round your answer to 3significant digits

OpenStudy (jfraser):

what size is the volumetric flask?

OpenStudy (anonymous):

450 mL

OpenStudy (anonymous):

sorry lol

OpenStudy (anonymous):

please help jfraser

OpenStudy (jfraser):

the units of concentration are \[M = \frac{mol \space solute}{L \space solution}\] You have enough information already to plug into the fraction and solve

OpenStudy (anonymous):

ok so out 25.7/450 ?

OpenStudy (anonymous):

im confused about where to place things

OpenStudy (anonymous):

i mean so put

OpenStudy (anonymous):

please help im confused

OpenStudy (jfraser):

the numbers are the least important thing to use, when you have to solve chemistry problems. Read the problem, and look at what the units say. 25.7 MICRO-moles in 450 mL but the units of molarity have to be MOLES per LITER. You have to pay attention to the prefixes that units sometimes come with.

OpenStudy (jfraser):

MICRO means 1 million times smaller, so 25.7 micromoles is 0.0000257 moles of KMnO4 MILLI means 1 thousand times smaller, so 450mL is 0.450L Those are the numbers that have to be plugged in, and where they go will be determined by the UNITS \[M = \frac{mol \space solute}{L \space solution} = \frac{0.0000257mol \space KMnO_4}{0.450L}\]

OpenStudy (anonymous):

so first convert the micro moles into moles

OpenStudy (anonymous):

and since the are asking for m/L places the new find moles over the size of the flask

OpenStudy (jfraser):

yes

OpenStudy (jfraser):

be careful using "m" for moles, since "m" is usually reserved for meters. Most everyone uses "mol" for moles, that way there's less confusion

OpenStudy (anonymous):

omg thank you so so much u saved me!

OpenStudy (anonymous):

im now your fan!

OpenStudy (jfraser):

anytime